Tutors please help me solve this equation:
Determine whether the ordered pairs are solutions of the equation
((((y=2/3x+3)))
a.3,5
b.-3,1
c.-6,0
d.0,3

Answer by THANApHD(104)

answer C is false , others are true
Y =2/3x+3
= 2/3*-6 + 3
= -4 + 3
= -1
So its false